If F =-kx and greater than 0, then what happens to the value of k if F triples in value and x remains constant
A)increase
B)doubles
C)decreases
D) cannot be determined

Answer by rothauserc(4718)

If you mean F > 0, then
:
F = -kx and F > 0 means that x < 0
:
F = kx
:
k = F/x
:
3F = kx
:
k = 3F/x
:
answer is A)increase
:
If you mean x > 0, then
:
F = -kx
:
-k = F/x
:
k = -F/x
:
3F = -kx
:
-k = 3F/x
:
k = -3F/x
:
answer is C)decreases
: